A boy got 50% of the questions on a test correct. If he had 10 questions correct out of the first 12, and 1/4 of the remaining questions correct, how many questions were on the test?

Answer by ewatrrr(24785)


Hi
Let x represent the number of questions on the test
Question states***
.50x = 10 + 1/4(x-12)
Solving for x
1/2x = 10 + 1/4x - 3
2x = 40 + x - 12
x = 28 questions on the test
CHECKING our Answer***
14 = 10 + 1/4(16)